OpenStudy (anonymous):

a particular fraction is equivalent to 2/3. the sum of its numerator and denominator is 105. what is the numerator of the fraction?

OpenStudy (raden):

just calculate ; 2/5 * 105

OpenStudy (lgbasallote):

let the particular fraction be x/y it says the fraction is equivalent to 2/3 so you can write it as x/y = 2/3 agree?

OpenStudy (raden):

yea, agree... i use smartway only

OpenStudy (lgbasallote):

i don't think that works though....

OpenStudy (anonymous):

agree but then the sum of the numerator and denominator is 10. would that be x+y=105 then?

OpenStudy (lgbasallote):

yes

OpenStudy (lgbasallote):

then from x/y = 2/3 you can cross multiply to get x = 2y/3 so substitute 2y/3 into x + y = 105

OpenStudy (raden):

my experience told me its same ^^

OpenStudy (lgbasallote):

we'll see

OpenStudy (raden):

yups..

OpenStudy (anonymous):

yes, it works. thanks!

OpenStudy (raden):

34, right ?

OpenStudy (raden):

ops, wrong... 42 :)